# Adventures in Demand Analysis Using AI

### v1

@@ 3.4 Orthogonal Inference of Causal Effects
[…]
@@
This claim follows directly from the regression equation (4), by defining
$e_{it}:=q_{t}\bigl{(}S_{it},\varepsilon_{it}\bigr{)}\;-\;\gamma_{t}\bigl{(}S_{it}\bigr{)},$
and then verifying the required orthogonality between $e_{it}$ and $P^{\perp}_{it}$.
Once we learn the CAPE, we effectively learn the CACE, provided that the causal SEM postulated above holds. If the SEM is only approximate, then the CAPE can still be treated as an approximation to the CACE; we elaborate on this in Section 3.6.
We estimate the CAPE and CACE using both linear regression and nonlinear, nonparametric models, applying modern machine-learning tools and cross-fitting to compute the residualized outcomes and exposure variables. We then estimate the projection equation (7) using either homogeneous or heterogeneous forms of $\delta_{t}(\cdot)=\alpha_{t}(\cdot)$ via least squares, and apply conventional statistical inference to construct $p$-values and confidence intervals, following Chernozhukov et al. (2018a). Details on the workflow and implementation algorithms are provided in the Online Appendix.
@@ 3.5.1 Homogeneous Elasticity Model
We begin by examining the homogeneous elasticity function. Table 7 shows that, in various specifications—ranging from simpler linear regressions with lagged price and quantity to more complex boosted-tree models incorporating product embeddings and additional controls—price consistently exerts a negative and highly significant effect on the quantity signal (the inverse sales rank). The confidence intervals range from about $-0.6$ to $-0.45$, indicating strong economic and statistical significance. Indeed, these estimates imply that a 1% price increase reduces the inverse sales rank by about $[0.6,0.45]\%$. Furthermore, to convert this effect into an elasticity of the actual demand under the power law assumption (cf. Remark 1), we need to multiply these coefficients by about 2, which indicates that the demand elasticity itself falls in the range of approximately $[-1.2,-0.9]$.
@@
Specification of Control Function (State $S_{t}$) coef std err t P-val. [5.0% 95.0%]
I-1. Linear ($P_{t-1}$, $Q_{t-1}$) -0.542 0.041 -13.372 $0$0.608 -0.475
I-1. Linear ($P_{t-1}$, $Q_{t-1}$, $X^{e},X^{o}_{t}$) -0.528 0.040 -13.208 $0$0.594 -0.463
I-1. Linear ($P_{t-1}$, $Q_{t-1}$, $X^{sim},X^{o}_{t}$) -0.527 0.040 -13.140 $0$0.593 -0.461
I-2. Linear with Interactions ($P_{t-1}$, $Q_{t-1}$, $X^{sim},X^{o}$) -0.529 0.040 -13.326 $0$0.595 -0.464
I-3. Boosted Trees ($P_{t-1}$, $Q_{t-1}$, $X^{e},X^{o}_{t}$) -0.538 0.038 -14.041 $0$0.601 -0.475
I-3. Boosted Trees ($P_{t-1}$, $Q_{t-1}$, $X^{sim},X^{o}_{t}$) -0.513 0.039 -13.191 $0$0.577 -0.449
@@ 3.5.2 Heterogeneous Elasticity Model
While the homogeneous specification yields more plausible estimates of average price sensitivity, it may still obscure important differences across products. The next step, therefore, models heterogeneous effects, allowing price elasticity to vary as a function of observed product characteristics or embedding-based similarity measures. Figure 6 and Table 8 illustrate how the elasticity estimates relate to these cluster similarities and to lagged price and quantity signals. These results reveal that certain products are more price-sensitive than others. For example, higher-priced and better-ranked products show greater price sensitivity, and similarities (to cluster centroids) also emerge as key drivers of sensitivity. (The first point is perhaps clearer from Table 8.) Products associated with specific clusters experience notably different price-sensitivity levels, underscoring the importance of accounting for product-level heterogeneity when estimating price effects.
@@
Modifier coef std err t P-val. [5.0% 95.0%]
Centercept -0.643 0.039 -16.643 $0$0.706 -0.579
Lagged Quantity, $Q_{t-1}$0.226 0.056 -4.035 $0$0.318 -0.134
Lagged Price, $P_{t-1}$0.167 0.031 -5.322 $0$0.219 -0.115
Cluster Similarity 00 -0.179 0.722 -0.248 0.804 -1.367 1.009
Cluster Similarity $1$7.134 2.890 -2.469 0.014 -11.887 -2.380
Cluster Similarity $2$ 0.630 0.499 1.261 0.207 -0.191 1.451
Cluster Similarity $3$6.932 2.905 -2.386 0.017 -11.710 -2.154
Cluster Similarity $4$ 2.615 1.158 2.259 0.024 0.711 4.519
@@
To provide a comprehensive view of how strong this heterogeneity can be, Figure 7 plots each product’s estimated elasticity (vertical axis) sorted in ascending order (horizontal axis is the “percentile index”); as per Chernozhukov et al. (2018b). The solid blue line represents the point estimates, the blue band denotes the 90% pointwise confidence intervals, and the dashed red line indicates the overall average elasticity. Reflecting the discussion of heterogeneous price sensitivity, elasticity ranges from about $-1.4$ to nearly 00, and these wide differences are statistically significant (as shown by the confidence bands). By multiplying these numbers by 2, we estimate that actual demand elasticity spans roughly $-2.8$ to 00, with an average near $-2.1$. This range highlights the importance of allowing for heterogeneity in price responses, as some products appear far more (or far less) sensitive to price changes than the average.
@@
II.1 Linear Specification 0.000 0.031
II.2 Linear Specification (Interactions) 0.000 0.031
II.3 PLR Specification (Boosted Trees) 0.000 0.077
Note: Standard errors are clustered at the product level. LR is estimated using OLS. The PLR model is estimated using DML with cross-fitted boosted trees (sample size: $24,895$).
3.6 Some Reflections on the Limitations of the Analysis
@@
However, in our setting, we suspect that the link $\epsilon_{it}\to P_{it}$ may be weak, as prices often follow “sticky,” piecewise-constant paths that do not change as frequently as quantity signals (sales ranks); see, for example, Figures 2 in Section 2. Formally, if the edge from $\epsilon_{it}$ to $P_{it}$ is zero, our earlier identification strategy holds, and our estimates are indeed causal. Otherwise, we can treat them as approximations of the causal estimates. For methods to bound the effect distortion caused when $\epsilon_{it}$ affects $P_{it}$, see Chernozhukov et al. (2021).
@@
Figure 8: Dynamic model with demand shocks as confounders.
Another way to achieve identification is through an instrumental variable $Z_{it}$ that induces exogenous variation in $P_{it}$, independent of $\epsilon_{it}$. By conditioning on the variation produced by $Z_{it}$, we isolate a component of $P_{it}$ uncorrelated with $\epsilon_{it}$, which then identifies the causal effect of price changes on quantity changes. Essentially, one can estimate the average causal effect of $Z_{it}$ on $Q_{it}$ (controlling for $S_{it}$), and the average causal effect of $Z_{it}$ on $\Delta P_{it}$ (also controlling for $S_{it}$), and then take their ratio—following the classical approach of Philip Wright (1928), republished as Wright (2024)—to identify the causal effect of $P_{it}$ on $Q_{it}$. We could not find such credible instruments in our setting. Using further lagged values of $Q_{i,t-1}$ is one potential approach, but it would require taking the first-order autoregressive specification for outcome too literally (i.e., those further lagged variables should likely be included as state variables, rather than employed as instruments).
@@ 4 Concluding Remarks
This study highlights the significant potential of AI-generated multimodal embeddings in demand analysis. By integrating text, image, and tabular data into a causal econometric framework, we improve both the precision of demand and price forecasts and the credibility of elasticity estimates. Our findings show that these rich embeddings not only enhance predictive accuracy but also reveal substantial heterogeneity in consumer price sensitivity, providing nuanced insights into demand behavior. These advancements create a methodological bridge between machine learning and econometrics, illustrating how modern AI tools can enrich traditional economic analyses. In future research, we hope to further explore the intersection of AI and causal inference, including bias-bounding and instrumental variable strategies to address time-varying confounding.
@@ Appendix A Addendum: Robustness to Estimated Embeddings
Let $Y_{it}$ denote the predictive target, which is either $P_{it}$ or $Q_{it}$. Let $X^{e}_{i}(\hat{\phi})$ denote the estimated and fine-tuned embeddings, where $\hat{\phi}$ denotes estimated parameters, and $S_{it}(\hat{\phi})$ the derived controls. We assume that $\hat{\phi}$ is obtained from data that are independent of the main data used in the analysis. Similarly, let $X_{i}^{e}$ and $S_{it}$ denote the ideal embeddings and controls, in the sense that
@@
In other words, after including $S_{it}$, the best prediction rule for $Y_{it}$ given both $S_{it}$ and $S_{it}(\hat{\phi})$ depends only on $S_{it}$.
Consider a learner $\hat{\gamma}^{Y}_{t}(S_{it}(\hat{\phi}))$ that minimizes empirical risk $\sum_{i\in A}\{Y_{it}-\gamma(S_{it}(\hat{\phi}))\}^{2}$ over control functions $\gamma$ in the convex model $\mathcal{F}$, conditional upon fine-tuned embeddings $\hat{\phi}$ (this includes our considered models 1-3). Here $A$ is a subset of $\{1,..,n\}$ whose size is proportional to $n$. The DML inference approaches using ideal and estimated embeddings are first-order equivalent under the following two key conditions:
@@
Work with the setup in this section. Assume that the data $W_{it}=(Q_{it},P_{it},X^{in}_{it})$, are identically distributed across $i$ for each $t=1,..,T$, where $T$ is fixed. Assume conditions (E1) and (E2) hold and that $Y_{it}$ and $\mathcal{F}$ are bounded. Then the empirical risk minimizer $\hat{\gamma}^{Y}_{t}(S_{it}(\hat{\phi}))$ learns $\mathrm{E}[Y_{it}\mid S_{it}]$ at the rate $o_{p}(n^{-1/4})$: $\max_{t}\sqrt{\mathrm{E}[\{\mathrm{E}[Y_{it}\mid S_{it}]-\hat{\gamma}(S_{it}(\hat{\phi}))\}^{2}\mid\hat{\phi}]}=o_{p}(n^{-1/2}).$
@@
First note that for any $\gamma\in\mathcal{F}$, the random variable $\hat{Z}_{it}(\gamma)\coloneqq\gamma(S_{it}(\hat{\phi}))$ belongs to the subspace $E\subset L^{2}$ consisting of functions that are measurable w.r.t. the pair $(S_{it},S_{it}(\hat{\phi}))$. Thus we have
@@
by the Pythagorean theorem, since $\mathrm{E}[Y_{it}|S_{it},S_{it}(\hat{\phi})]$ is the orthogonal projection of $Y_{it}$ onto $E$. It follows that the left-hand side depends on $\gamma$ only through: $\mathrm{E}\{\mathrm{E}[Y_{it}|S_{it},S_{it}(\hat{\phi})]-\hat{Z}_{it}(\gamma)\}^{2}=\mathrm{E}\{\mathrm{E}[Y_{it}|S_{it}]-\hat{Z}_{it}(\gamma)\}^{2}$, where we use the fact that $S_{it}$ is derived from the ideal embeddings.
@@
We describe the DML slope estimator next. Let $(I_{\ell})_{\ell}^{L}$ be the partition of $[n]=\{1,...,n\}$ into $L$ folds of approximately equal size. In step 1, for each $\ell$: obtain $\hat{\gamma}^{Y}_{t,\ell}(S_{it}(\hat{\phi}))$, the empirical risk minimizer over observation indices $A=[n]\setminus I_{\ell}$, for $Y=Q$ and $Y=P$; then obtain the residuals $\hat{P}_{it}^{\perp}=P_{it}-\hat{\gamma}^{P}_{t,\ell}(S_{it}(\hat{\phi}))$ and $\hat{Q}_{it}^{\perp}=Q_{it}-\hat{\gamma}^{Q}_{t,\ell}(S_{it}(\hat{\phi}))$ for $i\in I_{\ell}$. Then, obtain the slope estimator as:
@@
Work with conditions of the previous proposition. In addition, assume that $\mathrm{E}P_{it}^{\perp 2}$ is bounded away from zero. Then $\sqrt{n}(\hat{\delta}_{t}-\delta_{t})=(\mathrm{E}P^{\perp 2}_{it})^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}P^{\perp}_{it}e_{it}+o_{p}(1)\to_{d}N(0,V),$ where $V=\mathrm{E}[P^{\perp 2}_{it}]^{-2}\mathrm{E}[P^{\perp 2}_{it}e^{2}_{it}].$

### v2

@@ 3.4 Orthogonal Inference of Causal Effects
If Assumption 1(a) and (b) hold, then the CACE is identified by the CAPE:
@@
Proof.
Note that from equation (4) we have $Q_{it}=\alpha_{t}(S_{it})P_{it}+\gamma_{t}(S_{it})+u_{it}$, where $u_{it}:=q_{t}(S_{it},\varepsilon_{it})-\gamma_{t}(S_{it})$ satisfies $\mathrm{E}[u_{it}|S_{it}]=0$ by construction. Subtracting $\mathrm{E}[-|S_{it}]$ from both sides gives $Q_{it}^{\perp}=\alpha_{t}(S_{it})P_{it}^{\perp}+u_{it}$. Assumption 1 ensures that $\mathbb{E}[u_{it}P_{it}^{\perp}|S_{it}]=0$, since $u_{it}$ depends only on $(S_{it},\varepsilon_{it})$ while $P_{it}^{\perp}$ depends only on $(S_{it},\varepsilon_{it}^{p})$, and the error terms are mutually independent. Multiplying both sides by $P_{it}^{\perp}$ and taking conditional expectations w.r.t. $S_{it}$ then gives $\mathrm{E}[Q_{it}^{\perp}P_{it}^{\perp}|S_{i}t]=\alpha_{it}(S_{it})\mathrm{E}[(P_{it}^{\perp})^{2}|S_{it}]$, hence $\delta_{t}(S_{it})=\alpha_{t}(S_{it})$. ∎
It follows that once we learn the CAPE, we effectively learn the CACE, provided that the causal SEM postulated above holds. If the SEM is only approximate, then the CAPE can still be treated as an approximation to the CACE; we elaborate on this in Section 3.6.
We estimate the CAPE and CACE using both linear regression and nonlinear, nonparametric models, applying modern machine-learning tools and cross-fitting to compute the residualized outcomes and exposure variables. We then estimate the projection equation (7) using either homogeneous or heterogeneous forms of $\delta_{t}(\cdot)=\alpha_{t}(\cdot)$ via least squares, and apply conventional statistical inference to construct $p$-values and confidence intervals, following Chernozhukov et al. (2018a). Details on the workflow and the algorithms used are provided in the Online Appendix.
@@ 3.5.1 Homogeneous Elasticity Model
We begin by examining the homogeneous elasticity function. Table 7 shows that, in various specifications—ranging from simpler linear regressions with lagged price and quantity to more complex boosted-tree models incorporating product embeddings and additional controls—price consistently exerts a negative and highly significant effect on the quantity signal (the inverse sales rank). The confidence intervals range from about 0.79-0.79 to 0.54-0.54, indicating strong economic and statistical significance. Indeed, these estimates imply that a 1% price increase reduces the inverse sales rank by about $0.54,0.79]\%$. Furthermore, to convert this effect into an elasticity of the actual demand under the power law assumption (cf. Remark 1), we need to multiply these coefficients by about 2, which indicates that the demand elasticity itself falls in the range of approximately [1.58,1.08][-1.58,-1.08].
@@
Specification of Control Function (State $S_{t}$) coef std err t P-val. [5.0%, 95.0%]
I-1. Linear ($P_{t-1}$, $Q_{t-1}$) -0.690 0.040 -17.248 <0.001 -0.756 -0.624
I-1. Linear ($P_{t-1}$, $Q_{t-1}$, $X^{e},X^{o}_{t}$) -0.712 0.039 -18.364 <0.001 -0.776 -0.649
I-1. Linear ($P_{t-1}$, $Q_{t-1}$, $X^{sim},X^{o}_{t}$) -0.723 0.039 -18.725 <0.001 -0.786 -0.659
I-2. Linear with Interactions ($P_{t-1}$, $Q_{t-1}$, $X^{sim},X^{o}_{t}$) -0.727 0.039 -18.698 <0.001 -0.791 -0.662
I-3. Boosted Trees ($P_{t-1}$, $Q_{t-1}$, $X^{e},X^{o}_{t}$) -0.697 0.049 -14.362 <0.001 -0.704 -0.542
I-3. Boosted Trees ($P_{t-1}$, $Q_{t-1}$, $X^{sim},X^{o}_{t}$) -0.691 0.041 -17.051 <0.001 -0.777 -0.617
@@ 3.5.2 Heterogeneous Elasticity Model
While the homogeneous specification yields more plausible estimates of average price sensitivity, it may still obscure important differences across products. The next step, therefore, models heterogeneous effects, allowing price elasticity to vary as a function of observed product characteristics or embedding-based similarity measures. Figure 6 and Table 8 illustrate how the elasticity estimates relate to these cluster similarities and to lagged price and quantity signals. These results reveal that certain products are more price-sensitive than others. For example, higher-priced and better-ranked products show greater price sensitivity. [^2] This finding is most clearly seen in Table 8; note that the negative coefficient on lagged price lacks statistical significance, while the negative coefficient on lagged sales rank is highly significant. Similarities to cluster centroids also emerge as important drivers of sensitivity. Products associated with specific clusters experience notably different price-sensitivity levels, underscoring the importance of accounting for product-level heterogeneity when estimating price effects.
@@
Modifier coef std err t P-val. [5.0%, 95.0%]
Centercept -0.736 0.046 -15.860 <0.001 -0.812 -0.659
Lagged Quantity, $Q_{t-1}$ -0.325 0.092 -3.515 <0.001 -0.477 -0.173
Lagged Price, $P_{t-1}$ -0.049 0.035 -1.401 0.161 -0.106 0.008
Cluster Similarity 0 -2.312 1.524 -1.517 0.129 -4.819 0.195
Cluster Similarity 11 -7.365 1.831 -4.023 <0.001 -10.377 -4.354
Cluster Similarity 22 -6.785 3.558 -1.907 0.056 -12.637 -0.933
Cluster Similarity 33 -8.255 3.985 -2.072 0.038 -14.810 -1.700
Cluster Similarity 44 -4.559 1.303 -3.500 <0.001 -6.702 -2.417
@@
To provide a comprehensive view of how strong this heterogeneity can be, Figure 7 plots each product’s estimated elasticity (vertical axis) sorted in ascending order (horizontal axis is the “percentile index”); as per Chernozhukov et al. (2018b). The solid blue line represents the point estimates, the blue band denotes the 90% pointwise confidence intervals, and the dashed red line indicates the overall average elasticity. Reflecting the discussion of heterogeneous price sensitivity, elasticity ranges from about 2.0-2.0 to nearly $0.4$, and these wide differences are statistically significant (as shown by the confidence bands). By multiplying these numbers by 2, we estimate that actual demand elasticity spans roughly 4.0-4.0 to $0.8$, with an average near 1.5-1.5. This range highlights the importance of allowing for heterogeneity in price responses, as some products appear far more (or far less) sensitive to price changes than the average.
@@
II.1 Linear Specification <0.001 0.030
II.2 Linear Specification (Interactions) <0.001 0.054
II.3 PLR Specification (Boosted Trees) <0.001 0.001
Note: Standard errors are clustered at the product level. LR is estimated using OLS. The PLR model is estimated using DML with cross-fitted boosted trees (sample size: $38,041$).
3.6 Limitations of the Analysis and Sensitivity Bounds
@@
However, in our setting, we suspect that the link $\epsilon_{it}\to P_{it}$ may be weak, as prices often follow “sticky,” piecewise-constant paths that do not change as frequently as quantity signals (sales ranks); see, for example, Figure 2 in Section 2. Formally, if the edge from $\epsilon_{it}$ to $P_{it}$ is zero, our earlier identification strategy holds, and our estimates are indeed causal. Otherwise, we can treat them as approximations of the causal estimates.
For methods to bound the effect distortion caused when the omitted confounder $\epsilon_{it}$ affects $P_{it}$, we can employ the the senstitivity framework of Chernozhukov et al. (2021). Section 4 of the Online Appendix uses this framework to investigate how strong unobserved confounding would need to be to materially alter the estimated price effect. For concreteness, we focus on the nonlinear (boosted-trees) state-control estimate of the rank elasticity, -0.697 (Model I-3), and consider the adversarial case in which the omitted factor shifts demand and price in perfectly aligned directions so as to maximize bias. Allowing a latent confounder to explain up to 5% of the residual variation in both quantity and price yields an estimated rank-elasticity bound of [-0.919, -0.475], with corresponding $90\%$ one-side confidence bounds of -0.979 and -0.409, placing the effect squarely within the economically plausible range for rank elasticities. To obtain bounds wide enough to include zero while also accommodating very large elasticities in magnitude (near -1.5), the omitted confounder would need to explain roughly 15% of the residual variation in both outcome and price. This requirement is demanding relative to observed benchmarks: after adjusting for lagged outcomes and lagged prices, which proxy the relevant demand states, the remaining observed covariates contribute essentially no explanatory power for residual movements in either price or quantity, making the 5% adversarial-confounding benchmark already conservative.
@@
Figure 8: Dynamic model with demand shocks as omitted confounders.
Another way to achieve identification is through an instrumental variable $Z_{it}$ that induces exogenous variation in $P_{it}$, independent of $\epsilon_{it}$. By conditioning on the variation produced by $Z_{it}$, we isolate a component of $P_{it}$ uncorrelated with $\epsilon_{it}$, which then identifies the causal effect of price changes on quantity changes. Essentially, one can estimate the average causal effect of $Z_{it}$ on $Q_{it}$ (controlling for $S_{it}$), and the average causal effect of $Z_{it}$ on $\Delta P_{it}$ (also controlling for $S_{it}$), and then take their ratio—following the classical approach of Philip Wright (1928), republished as Wright (2024)—to identify the causal effect of $P_{it}$ on $Q_{it}$. We could not find such credible instruments in our setting. Using further lagged values of $Q_{i,t-1}$ is one potential approach, but it would require taking the first-order autoregressive specification too literally (i.e., those further lagged variables should likely be included as state variables, rather than employed as instruments).
@@ 4 Concluding Remarks
This study highlights the significant potential of AI-generated multimodal embeddings in demand analysis. By integrating text, image, and tabular data into a causal econometric framework, we improve both the precision of demand and price forecasts and the credibility of elasticity estimates. Our findings show that these rich embeddings not only enhance predictive accuracy but also reveal substantial heterogeneity in consumer price sensitivity, providing nuanced insights into demand behavior. These advancements create a methodological bridge between machine learning and econometrics, illustrating how modern AI tools can enrich traditional economic analyses.
Acknowledgments and Data Availability
We would like to thank the anonymous Associate Editor and a referee, and also Bulat Gafarov, Emi Nakamura, Patrick Kline, and Sylvia Klosin, for useful comments that improved the paper. We also thank the AI tools – Refine.Ink and ChatGPT-Pro – for additional helpful comments that we incorporated during the revision. Data and code are available at
https://anonymous.4open.science/r/demand-analysis-repro.
@@ Appendix A Addendum: Robustness to Estimated Embeddings
[…]
@@
[…]
Consider a learner $)\hat{\gamma}^{Y}_{t}(S_{it}(\hat{\phi}$)) that minimizes empirical risk $\sum_{i\in A}\{Y_{it}-\gamma(S_{it}(\hat{\phi}))\}^{2}$ over control functions $\gamma$ in the convex model $\mathcal{F}$, conditional upon fine-tuned embeddings $\hat{\phi}$ (this includes our considered models 1-3). Here $A$ is a subset of $\{1,..,n\}$ whose size is proportional to $n$. The DML inference approaches using ideal and estimated embeddings are first-order equivalent under the following two key conditions:
@@
Work with the setup in this section. Assume that the data $W_{it}=(Q_{it},P_{it},X^{in}_{it})$, are identically distributed across $i$ for each $t=1,..,T$, where $T$ is fixed. Assume conditions (E1) and (E2) hold and that $Y_{it}$ and $\mathcal{F}$ are bounded. Then the empirical risk minimizer $)\hat{\gamma}^{Y}_{t}(S_{it}(\hat{\phi}$)) learns $\mathrm{E}[Y_{it}\mid S_{it}]$ at the rate $o_{p}(n^{-1/4})$: $\max_{t}\sqrt{\mathrm{E}[\{\mathrm{E}[Y_{it}\mid S_{it}]-\hat{\gamma}(S_{it}(\hat{\phi}))\}^{2}\mid\hat{\phi}]}=o_{p}(n^{-1/4}).$
@@
First note that for any $\gamma\in\mathcal{F}$, the random variable $)\hat{Z}_{it}(\gamma)\coloneqq\gamma(S_{it}(\hat{\phi}$)) belongs to the subspace $E\subset L^{2}$ consisting of functions that are measurable w.r.t. the pair $)(S_{it},S_{it}(\hat{\phi}$)). Thus we have
@@
[…]
@@
We describe the DML slope estimator next. Let $(I_{\ell})_{\ell}^{L}$ be the partition of $n]=\{1,...,n\}$ into $L$ folds of approximately equal size. In step 1, for each $\ell$: obtain $)\hat{\gamma}^{Y}_{t,\ell}(S_{it}(\hat{\phi}$)), the empirical risk minimizer over observation indices $A=[n]\setminus I_{\ell}$, for $Y=Q$ and $Y=P$; then obtain the residuals $)\hat{P}_{it}^{\perp}=P_{it}-\hat{\gamma}^{P}_{t,\ell}(S_{it}(\hat{\phi}$)) and $)\hat{Q}_{it}^{\perp}=Q_{it}-\hat{\gamma}^{Q}_{t,\ell}(S_{it}(\hat{\phi}$)) for $i\in I_{\ell}$. Then, obtain the slope estimator as:
@@
Work with conditions of the previous proposition. In addition, assume that $\mathrm{E}[(P_{it}^{\perp})^{2}]$ is bounded away from zero. Then $\sqrt{n}(\hat{\delta}_{t}-\delta_{t})=\{\mathrm{E}[(P_{it}^{\perp})^{2}]\}^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}P^{\perp}_{it}e_{it}+o_{p}(1)\to_{d}N(0,V),$ where $V=\mathrm{E}[(P_{it}^{\perp})^{2}]^{-2}\mathrm{E}[(P_{it}^{\perp})^{2}e^{2}_{it}].$